OpenStudy (anonymous):

PLEASE HELP!!! MEDAL WILL BE REWARDED!!!!! The position of an object at time t is given by s(t) = -4 - 2t. Find the instantaneous velocity at t = 6 by finding the derivative.

OpenStudy (austinl):

This problem is just like the last one. We need to find the derivative of s(t), then plug in the given value for time.

OpenStudy (anonymous):

Okay, so would it be -8? t=-8? but that's not the velocity is it?

OpenStudy (anonymous):

oh no wait -16

OpenStudy (austinl):

\[s(t) = -4 - 2t\] Wait... did you copy this down correctly?

OpenStudy (anonymous):

Yes

OpenStudy (austinl):

Well, when you take the derivative of s(t), it is 0... are you sure?

OpenStudy (anonymous):

yes that is exactly how the question states it

OpenStudy (austinl):

Hmm... well is it multiple choice?

OpenStudy (anonymous):

no this one unfortunately is an essay question(do it yourself kinda problem)

OpenStudy (anonymous):

That's why im confused

OpenStudy (austinl):

Well, based on what I have seen... I would have to say the instantaneous velocity at t=6 has to be zero. \[s(t) = - 4 -2t\] \[s^{\prime}(t) = 0\] So, in the derivative there is no t to place the given value into...

OpenStudy (austinl):

Excuse me, \[s^{\prime}(t) = -2\]

OpenStudy (austinl):

So, -2 for an instantaneous velocity?

OpenStudy (anonymous):

okay, I will put that I guess, thank you so much for your help! also, do you know much about finding limits from looking a graph? Multiple choice...

OpenStudy (austinl):

The way that this problem is set up is kinda silly.

OpenStudy (austinl):

From a graph? Sure, I can take a crack at it. Post another question.

OpenStudy (anonymous):

Okay thanks I will now